A flask has a mass of 78.23 g when empty and 593.63 g when filled with water. When the same flask is filled with concentrated sulfuric acid, H2SO4, its mass is 1026.57 g. What is the density of concentrated sulfuric acid? (Assume water has a density of 1.00 g/cm3 at the temperature of the measurement.) 


1
Expert's answer
2021-02-03T07:15:55-0500

"m_f = 78.23 \;g \\

m_{fw} = 593.63 \;g \\

m_{water} = m_{fw} -m_f \\

= 593.63 -78.23 \\

= 515.4 \;g \\

V_{water} = V_{asid}=515.4 \;mL \\

m_{fa} = 1026.57 \;g \\

m_{acid} = m_{fa} -m_f \\

= 1026.57 -78.23 \\

= 948.34 \;g \\

ρ=\frac{m}{V} \\

ρ_{acid} = \frac{948.34}{515.4} \\

= 1.84 \;g/mL"

Answer: 1.84 g/mL